# A low-rank solver for conforming multipatch Isogeometric Analysis

### v3

@@ 1 Introduction
Isogeometric Analysis (IgA) has undergone significant development since the seminal paper [14]. Among other reasons, IgA is interesting as high-order PDEs solver. Indeed, IgA benefits from the approximation properties of splines, which, especially in the case of maximal regularity (splines of polynomial degree $p$ and continuity $C^{p-1}$) exhibits approximation properties superior to those of classical $C^{0}$ or discontinuous piecewise polynomials (see for example [2]). These approximation properties are very close to optimality in the sense of n-widths (see [10]).
@@
As with all high-order numerical methods, having efficient solvers is fundamental, and this has been a topic of research activity in the IgA community. Among the various solutions developed in this regard, low-rank compression techniques have been recently explored, leveraging the tensorial construction of isogeometric spaces at the level of each patch. See, for example, [20, 13, 16, 24], where the authors use a low-rank representation of the isogeometric Galerkin matrices. Low-rank tensor methods have been also exploited in the solution of IgA linear systems in [12, 21], based on the Tucker format, and in [3], that uses the tensor-trains approximation of the unknown. In these papers, a single tensorial patch is considered.
@@
In this work, we also consider tensorial patches, but in the multipatch case. This extension is not trivial since the multipatch setting is not globally tensorial. To circumvent this issue, instead of representing the unknown in the usual multipatch basis (obtained by a continuous gluing of the individual patch bases) we construct appropriate tensorial subdomains by pairing patches two by two, and represent the global unknown as the disjoint union of the unknowns of the subdomains. This representation is not unique, leading to a larger and singular linear system. However, the advantage is that, on each subdomain, we can use a solver that leverages tensorial properties, and construct then a block diagonal preconditioner for the global system. This is equivalent, in the full-rank case, to an overlapping Schwarz preconditioner for the system written in the standard basis, see [1]. An additional benefit is that the global unknown admits an approximation obtained from local low-rank approximations on the subdomains. This allows us to use the low-rank preconditioner from [21] on each subdomain. We then apply a truncated preconditioned conjugate gradient solver to our singular linear system.
@@
It is known that Krylov solvers perform well with singular systems, provided that the right-hand side lies in the range of the system matrix, see e.g. [17, 15, 25]. However, in our approach, low-rank compression is performed at each step, meaning that the iterates are not computed exactly: in this context, convergence is less understood. We also observe that truncated Krylov methods have only been introduced recently, and their theory is still under development, even for non-singular systems, see [18, 23, 27]. Although there are gaps in the mathematical understanding and therefore the convergence of the linear solver we use in this work is not guaranteed, our benchmarks show that the proposed method works well, similarly to the single-patch non-singular case considered in [21], and it delivers significant memory compression, up to two orders of magnitude compared to the full-rank case.
@@
We mention the preprint [4], which shares with the present paper the goal of tackling multipatch isogeometric problems by combining low-rank techniques with a domain decomposition approach. However, the approach presented in [4] is different from ours both in the choice of the tensor format (tensor trains), and in the domain decomposition paradigm (non-overlapping).
@@
The paper is organized as follows. In Section 2 we present the basics of IgA and of tensor calculus. The core of the paper is Section 3, where we propose the novel multipatch low-rank strategy. We present some numerical experiments in Section 4 while in the last section we draw some conclusions.
@@ 2.1 B-Splines
For B-splines properties, we refer to [5]. Multivariate B-splines are defined by tensor product of univariate B-splines. In this paper we focus on three-dimensional problems. Thus, given $m_{d}$ and $p_{d}$ positive integers, we introduce for $d=1,2,3$ the open knot vectors $\Xi_{d}:=\{\xi_{1,d},\dots,\xi_{m_{d}+p_{d}+1,d}\}$, the corresponding B-splines $\hat{b}_{d,i_{d}}$, $i_{d}=1,\ldots,m_{d}$, mesh-sizes $h_{d}$ and univariate spline spaces $\widehat{\mathcal{S}}^{p}_{h_{d}}$. The maximum of the mesh-sizes is denoted as $h:=\max\{h_{d}\ |\ d=1,2,3\}$. Note that for simplicity we are assuming that all univariate B-splines have the same degree. Multivariate B-splines $\widehat{B}_{\boldsymbol{i}}:[0,1]^{3}\rightarrow\mathbb{R}$ are defined as
@@ 2.2 Isogeometric spaces on a multipatch domain
$V_{ptc}^{(i)}:=\left\{\hat{v}_{h}\circ\mathcal{F}_{i}^{-1}\middle|\;\hat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(i)}\right\}.$
@@
We assume that the meshes are conforming at the patch interfaces, i.e. for all $i$ and $j$ such that $\partial\Omega^{(i)}\cap\partial\Omega^{(j)}\neq 0$ and the intersection is not a point, the interfaces are fully matching (see [14] for more details).
@@ 2.3 Model problem
${\varepsilon}(\underline{v}):=\frac{1}{2}\left(\nabla\underline{v}+(\nabla\underline{v})^{T}\right)$ is the symmetric gradient, while $\lambda$ and $\mu$ denote the material Lamé coefficients, that for simplicity we assume to be constant and positive and such that the Poisson ratio $\frac{\lambda}{2(\mu+\lambda)}$ is smaller than 0.5.
The corresponding discrete Galerkin problem that we want to solve is the following: find $\underline{u}_{h}\in[V_{h}]^{3}$ such that
@@ 2.4 The Tucker format
$=\sum_{r_{3}=1}^{R_{3}^{\mathbf{B}}R_{3}^{\mathbf{v}}}\sum_{r_{2}=1}^{R_{2}^{\mathbf{B}}R_{2}^{\mathbf{v}}}\sum_{r_{1}=1}^{R_{1}^{\mathbf{B}}R_{1}^{\mathbf{v}}}[\mathfrak{B}\otimes\mathfrak{C}]_{r_{1},r_{2},r_{3}}\left(B_{(3,r^{{}^{\prime}}_{3})}v_{(3,r^{{}^{\prime\prime}}_{3})}\right)\otimes\left(B_{(2,r^{{}^{\prime}}_{2})}v_{(2,r^{{}^{\prime\prime}}_{2})}\right)\otimes\left(B_{(1,r^{{}^{\prime}}_{1})}v_{(1,r^{{}^{\prime\prime}}_{1})}\right)$
where, in the last line, $r^{{}^{\prime}}_{d}$ and $r^{{}^{\prime\prime}}_{d}$ satisfy $r_{d}=r^{{}^{\prime\prime}}_{d}+\left(r^{{}^{\prime}}_{d}-1\right)n_{d}$, for $d=1,2,3$. Note that $\mathbf{B}\boldsymbol{v}$ is still a Tucker vector with multilinear rank $\left(R_{1}^{\mathbf{B}}R_{1}^{\mathbf{v}},R_{2}^{\mathbf{B}}R_{2}^{\mathbf{v}},R_{3}^{\mathbf{B}}R_{3}^{\mathbf{v}}\right)$.
@@ 3 Low-rank multipatch method
We introduce another subdivision of the computational domain, namely
$\Omega=\bigcup_{i=1}^{\mathcal{N}_{sub}}\Theta^{(i)}.$
Here, differently from the patches defined in Section 2.2, the subdomains $\Theta^{(j)}$ are allowed to overlap, i.e. the intersection of two different subdomains can have a non-empty interior.
@@
$V_{sub}^{(i)}=\left\{\hat{v}_{h}\circ\mathcal{G}_{i}^{-1}\middle|\;\hat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}\right\},$ (3.1)
@@
In the setting described in Section 2.2, the subdomains $\Theta^{(i)}$ and the corresponding spaces $V_{sub}^{(i)}$ satisfying (3.1) can be obtained by merging pairs of isogeometric patches and the corresponding spaces We give details for this construction in Section 3.4.
@@ 3.1 The linear system
The functions $\underline{u}_{h}^{(1)},\ldots,\underline{u}_{h}^{(\mathcal{N}_{sub})}$ represent the unknowns. Due to the overlap between the subdomains, the subspaces $\left[V^{(1)}_{sub}\right]^{3},\ldots,\left[V^{(\mathcal{N}_{sub})}_{h}\right]^{3}$ are not necessarily in direct sum. As a result, $\underline{u}_{h}^{(1)},\ldots,\underline{u}_{h}^{(\mathcal{N}_{sub})}$ are not uniquely determined in general. Instead, they are selected by the Krylov iterative solver, based on the initial iterate, without the need to impose any average condition.
@@
$\mathbf{A}^{(i,j,k,\ell)}\approx\widetilde{\mathbf{A}}^{(i,j,k,\ell)}=\sum_{r_{3}=1}^{R_{3}^{A}}\sum_{r_{2}=1}^{R_{2}^{A}}\sum_{r_{1}=1}^{R_{1}^{A}}\mathfrak{A}_{r_{1},r_{2},r_{3}}\widetilde{A}^{(i,j,k,\ell)}_{(3,r_{3})}\otimes\widetilde{A}^{(i,j,k,\ell)}_{(2,r_{2})}\otimes\widetilde{A}^{(i,j,k,\ell)}_{(1,r_{1})},\qquad$ (3.5)
$\boldsymbol{f}^{(i,k)}\approx\widetilde{\boldsymbol{f}}^{(i,k)}=\sum_{r_{3}=1}^{R_{3}^{f}}\sum_{r_{2}=1}^{R_{2}^{f}}\sum_{r_{1}=1}^{R_{1}^{f}}\mathfrak{f}_{r_{1},r_{2},r_{3}}\tilde{f}^{(i,k)}_{(3,r_{3})}\otimes f^{(i,k)}_{(2,r_{2})}\otimes f^{(i,k)}_{(1,r_{1})},$ (3.6)
for $i,j=1,\ldots,\mathcal{N}_{sub}$, $k,\ell=1,2,3$, where $R^{A}_{d}=R^{A}_{d}(i,j,k,\ell)$ and $R^{f}_{d}=R^{f}_{m}(i,k)$ for $d=1,2,3$. In Section 3.4 we give details on how these approximations can be performed, when the subdomains are defined as unions of neighboring patches. We emphasize that, for the proposed strategy to be beneficial, the multilinear ranks of the approximations, which depend on the geometry coefficients and on the forcing term, have to be reasonably small.
@@
$\widetilde{\mathbf{A}}\widetilde{\boldsymbol{u}}=\widetilde{\boldsymbol{f}}.$
@@ 3.2 The preconditioner
We recall that, as discussed in [1, Theorem 2], the use of a block diagonal preconditioner is equivalent to the use of additive overlapping Schwarz preconditioners for the nonsingular linear system obtained by constructing a global spline basis. Here, however, each diagonal block $\left(\mathbf{P}^{(i,k)}\right)^{-1}$ is further approximated with a low-rank Tucker matrix following the steps described in [21, Section 4.2]. We emphasize in particular that this strategy is not spoiled by the presence of the constant coefficients. A key feature of this approach is that the action of $\left(\widetilde{\mathbf{P}}^{(i,k)}\right)^{-1}$ on a vector can be computed using a variant of the fast diagonalization method [19, 26] that exploits the fast Fourier transform (FFT), yielding almost linear complexity. In conclusion, the preconditioner $\widetilde{\mathbf{P}}$ is the block diagonal matrix obtained by collecting all matrices $\widetilde{\mathbf{P}}^{(i,k)}$, for $i=1,\ldots,\mathcal{N}_{sub}$, $k=1,2,3$.
@@
$\boldsymbol{v}^{T}\mathbf{A}^{(i,i,k,k)}\boldsymbol{v}=\int_{\Theta^{(i)}}2\mu\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{k}v_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\mathrm{d}{\underline{x}}=\int_{[0,1]^{3}}\left(\nabla\hat{v}_{sub}^{(i)}\right)^{T}Q^{(i,k)}\nabla\hat{v}_{sub}^{(i)}\mathrm{d}{\underline{\xi}},$
where $\hat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}\in\widehat{\boldsymbol{\mathcal{S}}}^{(i)}_{sub,0}$, and
@@
$\boldsymbol{v}^{T}\mathbf{P}^{(i,k)}\boldsymbol{v}=\int_{[0,1]^{3}}\left(\nabla\hat{v}_{sub}^{(i)}\right)^{T}\begin{bmatrix}c^{(i,k)}_{1}&0&0\\ 0&c^{(i,k)}_{2}&0\\ 0&0&c^{(i,k)}_{3}\end{bmatrix}\nabla\hat{v}_{sub}^{(i)}\mathrm{d}{\underline{\xi}}.$

### v4

@@ 1 Introduction
Isogeometric Analysis (IgA) has seen substantial advancements since the publication of the seminal work [22]. Among its many advantages, IgA stands out as an effective and versatile high-order method for approximating partial differential equations (PDEs). Indeed, IgA benefits from the approximation properties of splines, which, especially in the case of maximal regularity (splines of polynomial degree $p$ and continuity $C^{p-1}$) exhibits approximation properties superior to those of classical $C^{0}$ or discontinuous piecewise polynomials (see for example [8] and [18]).
@@
As with all high-order numerical methods, the development of efficient solvers is crucial, and this has been an active area of research within the IgA community. Among the various approaches proposed, low-rank compression techniques have recently garnered attention. These methods take advantage of the tensorial structure inherent in isogeometric spaces at the patch level. See, for example, [32, 21, 24, 39], where the authors use a low-rank representation of the isogeometric Galerkin matrices. Low-rank tensor methods have been also exploited in the solution of IgA linear systems in [20, 34], based on the Tucker format, and in [9], that uses the tensor-trains approximation of the unknown. In all these papers, a single tensorial patch is considered. This is a limitation that we aim to overcome.
@@
In this work, we indeed consider a multipatch domain. This extension is non-trivial since the multipatch setting is generally not globally tensorial, even when the individual patches are, as we assume. The use of classical domain decomposition techniques allows leveraging tensoriality at the level of local problems but not at the global level, see our previous work [6]. To circumvent this issue, in the present work, Instead of representing the unknown in the usual multipatch basis (obtained by a continuous gluing of the individual patch bases) we construct appropriate tensorial subdomains by merging neighboring patches, and represent the global unknown as the disjoint union of the unknowns of the subdomains. This representation is not unique, leading to a singular linear system. However, the advantage is that, on each subdomain, we can use a solver that leverages tensorial properties, and construct then a block diagonal preconditioner for the global system. This is equivalent, in the full-rank case, to an overlapping Schwarz preconditioner for the system written in the standard basis, see [7]. An additional benefit is that the global unknown admits a low-rank approximation on each subdomain. This allows us to use the low-rank preconditioner from [34] on each subdomain and, a truncated preconditioned conjugate gradient solver to solve our global singular linear system.
@@
It is known that Krylov solvers perform well with singular systems, provided that the right-hand side lies in the range of the system matrix, see e.g. [25, 23, 40]. However, in our approach, low-rank compression is performed at each step, meaning that the iterates are not computed exactly: in this context, convergence is less understood. We also observe that truncated Krylov methods have only been introduced recently, and their theory is still under development, even for non-singular systems, see [29, 38, 42]. Although there are gaps in the mathematical analysis and therefore the convergence of the linear solver we use in this work is not proved, our benchmarks show that the proposed method works well, similarly to the single-patch non-singular case considered in [34]. Moreover, it achieves substantial memory compression, reducing storage requirements by up to two orders of magnitude compared to the full-rank case.
@@
In the context of Isogeometric Analysis, only one work, the preprint [10], shares the objective of this paper: addressing multipatch isogeometric problems by combining low-rank techniques with a domain decomposition approach. However, it employs a different tensor format (tensor trains) and a non-overlapping domain decomposition method in a simplified setting.
Beyond Isogeometric Analysis, there is significant research activity focused on the use of low-rank methods for solving PDEs. We refer to the recent survey [3] for a throughout discussion on this topic. In particular, several low-rank solvers have been proposed in the literature, including Krylov subspace methods with rank truncation [26, 29, 4, 5, 30, 1], alternating direction approaches [35, 16, 36], Riemannian optimization [28, 44], and greedy strategies [2, 11]. The target problems of these works are often parametrized or stochastic PDEs, and the focus is on high dimension. We emphasize, however, that in all these references PDEs are solved on simple tensor-product domains. To the best of our knowledge, the only work proposing a low-rank solver for complex domains is [33] (and its follow-up [27]), but this approach, different from the one presented here, is restricted to 2-dimensional problems.
@@
The paper is organized as follows. In Section 2 we present the basics of IgA and of tensor calculus. The core of the paper is Section 3, where we propose the novel multipatch low-rank strategy. In Section 4 we present a theoretical study on the error committed on the system matrix due to the low-rank approximation of the geometry. We present some numerical experiments in Section 5 while in the last section we draw some conclusions.
@@ 2.1 B-Splines
For B-splines properties, we refer to [12]. Multivariate B-splines are defined by tensor product of univariate B-splines. In this paper we focus on three-dimensional problems. Thus, given $m_{d}$ and $p_{d}$ positive integers, we introduce for $d=1,2,3$ the open knot vectors $\Xi_{d}:=\{\xi_{1,d},\dots,\xi_{m_{d}+p_{d}+1,d}\}$, the corresponding B-splines $\widehat{b}_{d,i_{d}}$, $i_{d}=1,\ldots,m_{d}$, mesh-sizes $h_{d}$ and univariate spline spaces $\widehat{\mathcal{S}}^{p}_{h_{d}}$. The maximum of the mesh-sizes is denoted as $h:=\max\{h_{d}\ |\ d=1,2,3\}$. Note that for simplicity we are assuming that all univariate B-splines have the same degree. Multivariate B-splines $\widehat{B}_{\boldsymbol{i}}:[0,1]^{3}\rightarrow\mathbb{R}$ are defined as
@@ 2.2 Isogeometric spaces on a multipatch domain
$V_{ptc}^{(i)}:=\left\{\widehat{v}_{h}\circ\mathcal{F}_{i}^{-1}\middle|\;\widehat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{ptc,0}^{(i)}\right\}.$
@@
We assume that the intersection between two patches, when it is not empty, is either a full face, a full edge or a vertex. Moreover, we require that the meshes are conforming at the patch interfaces, i.e. for all $i$ and $j$ such that $\partial\Omega^{(i)}\cap\partial\Omega^{(j)}\neq 0$ and the intersection is not a point.
@@ 2.3 Model problem
Here ${\varepsilon_{i,j}}(\underline{v})=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)$ for $i,j=1,2,3$, while $\lambda$ and $\mu$ denote the material Lamé coefficients, that for simplicity we assume to be constant and positive and such that the Poisson ratio $\frac{\lambda}{2(\mu+\lambda)}$ is smaller than 0.5. The corresponding discrete Galerkin problem that we want to solve is the following: find $\underline{u}_{h}\in[V_{h}]^{3}$ such that
@@ 2.4 The Tucker format
$=\sum_{r_{3}=1}^{R_{3}^{\mathbf{B}}R_{3}^{\boldsymbol{v}}}\sum_{r_{2}=1}^{R_{2}^{\mathbf{B}}R_{2}^{\boldsymbol{v}}}\sum_{r_{1}=1}^{R_{1}^{\mathbf{B}}R_{1}^{\boldsymbol{v}}}[\mathfrak{B}\otimes\mathfrak{C}]_{r_{1},r_{2},r_{3}}\left(B_{(3,r^{{}^{\prime}}_{3})}v_{(3,r^{{}^{\prime\prime}}_{3})}\right)\otimes\left(B_{(2,r^{{}^{\prime}}_{2})}v_{(2,r^{{}^{\prime\prime}}_{2})}\right)\otimes\left(B_{(1,r^{{}^{\prime}}_{1})}v_{(1,r^{{}^{\prime\prime}}_{1})}\right)$
where, in the last line, $r^{{}^{\prime}}_{d}$ and $r^{{}^{\prime\prime}}_{d}$ satisfy $r_{d}=r^{{}^{\prime\prime}}_{d}+\left(r^{{}^{\prime}}_{d}-1\right)n_{d}$, for $d=1,2,3$. Note that $\mathbf{B}\boldsymbol{v}$ is still a Tucker vector with multilinear rank $\left(R_{1}^{\mathbf{B}}R_{1}^{\boldsymbol{v}},R_{2}^{\mathbf{B}}R_{2}^{\boldsymbol{v}},R_{3}^{\mathbf{B}}R_{3}^{\boldsymbol{v}}\right)$.
@@ 3 Low-rank multipatch method
We introduce another subdivision of the computational domain, namely let $\Theta^{(1)},\ldots,\Theta^{(\mathcal{N}_{sub})}$ be closed subsets of $\Omega$ such that
$\text{int}\left(\Omega\right)=\bigcup_{i=1}^{\mathcal{N}_{sub}}\text{int}\left(\Theta^{(i)}\right).$ (3.1)
where $\text{int}\left(\cdot\right)$ denotes the interior of a set. Here, differently from the patches defined in Section 2.2, the subdomains $\Theta^{(j)}$ are allowed to overlap, i.e. the intersection of two different subdomains can have a non-empty interior.
@@
Note that (3.1) guarantees that $V_{h}=\bigcup_{i=1}^{\mathcal{N}_{sub}}V^{(i)}_{sub}.$
@@
$V_{sub}^{(i)}=\left\{\widehat{v}_{h}\circ\mathcal{G}_{i}^{-1}\middle|\;\widehat{v}_{h}\in\widehat{\boldsymbol{\mathcal{S}}}_{sub,0}^{(i)}\right\},$ (3.2)
@@
In the setting described in Section 2.2, the subdomains $\Theta^{(i)}$ and the corresponding spaces $V_{sub}^{(i)}$ satisfying (3.2) can be obtained by merging isogeometric patches as well as the corresponding spaces. We give details for this construction in Section 3.4.
@@ 3.1 The linear system
The functions $\underline{u}_{h}^{(1)},\ldots,\underline{u}_{h}^{(\mathcal{N}_{sub})}$ represent the unknowns. Due to the overlap between the subdomains, the subspaces $\left[V^{(1)}_{sub}\right]^{3},\ldots,\left[V^{(\mathcal{N}_{sub})}_{sub}\right]^{3}$ are not necessarily in direct sum. As a result, $\underline{u}_{h}^{(1)},\ldots,\underline{u}_{h}^{(\mathcal{N}_{sub})}$ are not uniquely determined in general. Instead, they are selected by the Krylov iterative solver, based on the initial iterate, without the need to impose any average condition.
@@
$\mathbf{A}^{(i,j,k,\ell)}\approx\widetilde{\mathbf{A}}^{(i,j,k,\ell)}=\sum_{r_{3}=1}^{R_{3}^{\mathbf{A}}}\sum_{r_{2}=1}^{R_{2}^{\mathbf{A}}}\sum_{r_{1}=1}^{R_{1}^{\mathbf{A}}}\mathfrak{A}_{r_{1},r_{2},r_{3}}\widetilde{A}^{(i,j,k,\ell)}_{(3,r_{3})}\otimes\widetilde{A}^{(i,j,k,\ell)}_{(2,r_{2})}\otimes\widetilde{A}^{(i,j,k,\ell)}_{(1,r_{1})},\qquad$ (3.6)
$\boldsymbol{f}^{(i,k)}\approx\widetilde{\boldsymbol{f}}^{(i,k)}=\sum_{r_{3}=1}^{R_{3}^{\boldsymbol{f}}}\sum_{r_{2}=1}^{R_{2}^{\boldsymbol{f}}}\sum_{r_{1}=1}^{R_{1}^{\boldsymbol{f}}}\mathfrak{f}_{r_{1},r_{2},r_{3}}\widetilde{f}^{(i,k)}_{(3,r_{3})}\otimes f^{(i,k)}_{(2,r_{2})}\otimes f^{(i,k)}_{(1,r_{1})},$ (3.7)
for $i,j=1,\ldots,\mathcal{N}_{sub}$, $k,\ell=1,2,3$, where $R^{\mathbf{A}}_{d}=R^{\mathbf{A}}_{d}(i,j,k,\ell)$ and $R^{\boldsymbol{f}}_{d}=R^{\boldsymbol{f}}_{d}(i,k)$ for $d=1,2,3$. In Section 3.4 we give details on how these approximations can be performed, when the subdomains are defined as unions of neighboring patches. We emphasize that, for the proposed strategy to be beneficial, the multilinear ranks of the approximations, which depend on the geometry coefficients and on the forcing term, have to be reasonably small.
@@
$\widetilde{\mathbf{A}}\widetilde{\boldsymbol{u}}=\widetilde{\boldsymbol{f}}.$ (3.8)
@@ 3.2 The preconditioner
We recall that, as discussed in [7, Theorem 2], the use of a block diagonal preconditioner is equivalent to the use of additive overlapping Schwarz preconditioners for the nonsingular linear system obtained by constructing a global spline basis. Here, however, each diagonal block $\left(\mathbf{P}^{(i,k)}\right)^{-1}$ is further approximated with a low-rank Tucker matrix following the steps described in [34, Section 4.2]. We emphasize in particular that this strategy is not spoiled by the presence of the constant coefficients. A key feature of this approach is that the action of $\left(\widetilde{\mathbf{P}}^{(i,k)}\right)^{-1}$ on a vector can be computed using a variant of the fast diagonalization method [31, 41] that exploits the fast Fourier transform (FFT), yielding almost linear complexity. In conclusion, the preconditioner $\widetilde{\mathbf{P}}$ is the block diagonal matrix obtained by collecting all matrices $\widetilde{\mathbf{P}}^{(i,k)}$, for $i=1,\ldots,\mathcal{N}_{sub}$, $k=1,2,3$.
@@
$\boldsymbol{v}^{T}\mathbf{A}^{(i,i,k,k)}\boldsymbol{v}=\int_{\Theta^{(i)}}2\mu\varepsilon(\underline{e}_{k}v_{h}):\varepsilon(\underline{e}_{k}v_{h})+\lambda\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\left(\nabla\cdot\underline{e}_{k}v_{h}\right)\mathrm{d}{\underline{x}}=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{sub}^{(i)}\right)^{T}Q^{(i,k)}\nabla\widehat{v}_{sub}^{(i)}\mathrm{d}{\underline{\xi}},$
where $\widehat{v}_{sub}^{(i)}=v_{h}\circ\mathcal{G}_{i}\in\widehat{\boldsymbol{\mathcal{S}}}^{(i)}_{sub,0}$, and
@@
$\boldsymbol{v}^{T}\mathbf{P}^{(i,k)}\boldsymbol{v}=\int_{[0,1]^{3}}\left(\nabla\widehat{v}_{sub}^{(i)}\right)^{T}\begin{bmatrix}c^{(i,k)}_{1}&0&0\\ 0&c^{(i,k)}_{2}&0\\ 0&0&c^{(i,k)}_{3}\end{bmatrix}\nabla\widehat{v}_{sub}^{(i)}\mathrm{d}{\underline{\xi}}.$